A company estimates that 0.3% of their products will fail after the original warranty period but within 2 years of the purchase, with a replacement cost of $350.
If they offer a 2 year extended warranty for $29, what is the company's expected value of each warranty sold?
$

Answer by ikleyn(52863)

.
A company estimates that 0.3% of their products will fail after the original warranty period
but within 2 years of the purchase, with a replacement cost of $350.
If they offer a 2 year extended warranty for $29, what is the company's expected value of each warranty sold?

Solution

Notice that 0.3% = 0.003.


The expected profit of selling each warranty is

    29 - 0.003*350  dollars = 29 - 1.05 dollars = 27.95  dollars.    ANSWER


Here $29 is what a customer pays for the 2-years warranty,
and 0.003*350 = 1.05 dollars is the company's average expense per unit to cover the replacement.


The net  29 - 1.05 = 27.95 dollars is the expected company's profit of each warranty sold.


$27.95 is the price which an average customer does agree to pay to the company for peace in mind.

Solved and explained.